# biblio.bib

### v1

@@ A.2. Limit for a sequence of exponentially growing balls
We now follow [LieLeb-72, LieNar-75] and introduce
@@
Let $B_{k}:=B(0,R_{k})$. By [LieLeb-72, Lemma 4.1], one can pack the ball $B_{K}$ by a disjoint union of $n_{K-j}$ translated copies of the $B_{j}$’s with $j=0,...,K-1$, in such a way that the mutual distance between all the balls, including $\partial B_{K}$, is at least equal to $\frac{p}{2(1+p)}$ (Figure 3). A simple calculation shows that the total occupied volume satisfies
@@
Figure 3. Packing the large ball $B_{K}$ with smaller balls $B_{j}$ as in [LieLeb-72, LieNar-75] so that the uncovered part (in yellow in the picture) has an exponentially small fraction of the total volume $|B_{K}|$.
Let $\Omega_{j}$ be a minimizer for the problem $E_{B_{j}}(\rho)$ in the ball $B_{j}=B(0,R_{j})$. We call $x_{j,n}$ the centers of the packing balls of radii $R_{j}$ that are placed in the large ball $B_{K}$. In each of these balls, we place the trial set $\Omega_{j,n}:=x_{j,n}+U_{j,n}\Omega_{j}$ with a rotation $U_{j,n}\in SO(3)$. When we average over $U_{j,n}$ the Coulomb potential induced by ${\mathds{1}}_{\Omega_{j,n}}-\rho{\mathds{1}}_{B(x_{j,n},R_{j})}$ vanishes outside of the corresponding ball $B(x_{j,n},R_{j})$ by Newton’s theorem. This is due to the neutrality condition $|\Omega_{j}|=\rho|B_{j}|$. The Coulomb interaction between this ball and the rest of the system thus vanishes. We do this for all the small balls.
@@
Like in [LieNar-75] (and unlike in [LieLeb-72]), we cannot leave $D_{K}$ empty because the Coulomb energy of the background $\rho{\mathds{1}}_{D_{K}}$ is too large. Whatever $\Omega$ we put in $D_{K}$ will not interact with the balls, since those generate a Coulomb potential that exactly vanishes outside of the balls, as we have already explained. Hence after taking a minimizing $\Omega\subset D_{K}$ we obtain the bound
@@
exists. This is a consequence of the following lemma, which is essentially contained in [LieLeb-72, p. 343] (see also [LieSei-09, p. 264]) and whose proof we give for the convenience of the reader.
@@
Let $f_{j}$ be a sequence of real numbers that is bounded from below. Let $10<\gamma<$ 1 and assume that there exists a summable sequence $\delta_{j}\geqslant 0$ such that
@@
If therefore only remains to explain how to show (42). The construction of a competitor $\Omega$ for $E_{D_{K}}(\rho)$ is the main difficulty, compared with what was done in [LieNar-75]. It is harder to screen the background in the liquid drop model because ${\mathds{1}}_{\Omega}$ has a fixed height. Our idea is to split $D_{K}$ into many small boxes of size $\varepsilon$ and place in each full cube a smaller cube of side length $\rho^{1/3}\varepsilon$. Of course the main difficulty is to deal with the cubes intersecting the boundary of $D_{K}$ and this is where Proposition 10 becomes useful.
@@
Since the $\mu_{\alpha}$ are neutral, have no dipole and are contained in a set of diameter less that $C\varepsilon$, we have by  [AnaLew-20, Lemma 2.1]
@@ A.3. Upper bound for any regular set
Next we turn to the upper bound for an arbitrary sequence of sets $\Lambda_{n}\nearrow\mathbb{R}^{3}$ satisfying the conditions of Theorem 1. The idea is to again pack $\Lambda_{n}$ with many balls congruent to $B_{0},...,B_{K-1}$ for some fixed $K$. To this end, it is easier to start packing the large cube $C^{\prime}_{0}=|B_{K}|^{1/3}(-1/2,1/2)^{3}$ of the same volume as $B_{K}$ and then use this cube to pack $\Lambda_{n}$. By [LieLeb-72, Lemma 4.1] the cube $C^{\prime}_{0}$ can be packed by the same number $n_{K-j}$ of copies of the ball $B_{j}$, with the same estimate on the distance between the boundaries.
@@
#$\#\mathcal{Z}^{\prime}_{b}\leqslant|B_{K}|^{-1}\left|\left\{x\in\Lambda_{n}\ :\ \mathrm{d}(x,\partial\Lambda_{n})\leqslant(\sqrt{3}+1)h_{K}\right\}\right|\leqslant C\ell_{n}^{2}h_{K}$
@@
#$\#\mathcal{Z}^{\prime}_{i}=|B_{K}|^{-1}|\Lambda_{n}|+O(\ell_{n}^{2}h_{K}).$
@@
#$\leqslant\#\mathcal{Z}^{\prime}_{i}\sum_{j=0}^{K-1}n_{K-j}E_{B_{j}}(\rho)+C\gamma^{K}|\Lambda_{n}|+C\ell_{n}^{2}h_{K}.$ (44)
@@
Figure 4. Idea of the proof of the thermodynamic limit for an arbitrary sequence $\Lambda_{n}\nearrow\mathbb{R}^{3}$, following [LieLeb-72, LieNar-75]. To get an upper bound on $E_{\Lambda_{n}}(\rho)$ we pack the domain $\Lambda_{n}$ by many small balls, using a cubic tiling as guiding principle (only the largest balls of the type $B_{K-1}$ are displayed in the picture). To get a lower bound on $E_{\Lambda_{n}}(\rho)$ we pack the complement of $\Lambda_{n}$ in a large ball $B_{K_{n}}$ of comparable volume.
@@
Here $\mathcal{Z}^{\prime}_{i}$ denotes the sets of indices $z\in\mathbb{Z}^{3}$ such that $C^{\prime}_{z}$ is included in $B_{K_{n}}\setminus\Lambda_{n}$ at distance $>h_{K}$ from its boundary.
@@
The next step is to handle the Coulomb interaction between the quadrupoles and $\Lambda_{n}$ in (45). We use that the Coulomb potential generated by a quadrupole decays like $1/|x|^{3}$ [AnaLew-20, Lemma 2.1], that is,
@@ A.5. The periodic problem
That the periodic model admits the same limit $e(\rho)$, as stated in (7), can be shown by following [LewLieSei-19b, Section V]. Let us call
@@
To obtain the reverse bound, we pack $\Lambda_{L}$ with balls as explained in Subsection A.3 and in [LewLieSei-19b]. We choose $\varepsilon\in L/\mathbb{N}$ to avoid having any partial cubes at the boundary of $\Lambda_{L}$ (and thus be consistent with the periodic boundary condition). Using that Newton’s theorem applies in the periodic case as well, we find the same bound as (44) on $E_{L}^{\rm per}(\rho)$ and this gives

### v2

@@ A.2. Limit for a sequence of exponentially growing balls
We now follow [LL72, LN75] and introduce
@@
Let $B_{k}:=B(0,R_{k})$. By [LL72, Lemma 4.1], one can pack the ball $B_{K}$ by a disjoint union of $n_{K-j}$ translated copies of the $B_{j}$’s with j=0,,K1j=0,...,K-1, in such a way that the mutual distance between all the balls, including $\partial B_{K}$, is at least equal to $\frac{p}{2(1+p)}$ (Figure 3). A simple calculation shows that the total occupied volume satisfies
@@
Figure 3. Packing the large ball $B_{K}$ with smaller balls $B_{j}$ as in [LL72, LN75] so that the uncovered part (in yellow in the picture) has an exponentially small fraction of the total volume $|B_{K}|$.
Let $\Omega_{j}$ be a minimizer for the problem $E_{B_{j}}(\rho)$ in the ball $B_{j}=B(0,R_{j})$. We call $x_{j,n}$ the centers of the packing balls of radii $R_{j}$ that are placed in the large ball $B_{K}$. In each of these balls, we place the trial set $\Omega_{j,n}:=x_{j,n}+U_{j,n}\Omega_{j}$ with a rotation $U_{j,n}\in SO(3)$. When we average over $U_{j,n}$ the Coulomb potential induced by ${\mathds{1}}_{\Omega_{j,n}}-\rho{\mathds{1}}_{B(x_{j,n},R_{j})}$ vanishes outside of the corresponding ball $B(x_{j,n},R_{j})$ by Newton’s theorem as in (23). This is due to the neutrality condition $|\Omega_{j}|=\rho|B_{j}|$. The Coulomb interaction between this ball and the rest of the system thus vanishes. We do this for all the small balls.
@@
Like in [LN75] (and unlike in [LL72]), we cannot leave $D_{K}$ empty because the Coulomb energy of the background $\rho{\mathds{1}}_{D_{K}}$ is too large. Whatever $\Omega$ we put in $D_{K}$ will not interact with the balls, since those generate a Coulomb potential that exactly vanishes outside of the balls, as we have already explained. Hence after taking a minimizing $\Omega\subset D_{K}$ we obtain the bound
@@
exists. This is a consequence of the following lemma, which is essentially contained in [LL72, p. 343] (see also [LS09, p. 264]) and whose proof we give for the convenience of the reader.
@@
Let $f_{j}$ be a sequence of real numbers that is bounded from below. Let $0<\gamma<1$ and assume that there exists a summable sequence $\delta_{j}\geqslant 0$ such that
@@
If therefore only remains to explain how to show (44). The construction of a competitor $\Omega$ for $E_{D_{K}}(\rho)$ is the main difficulty, compared with what was done in [LN75]. It is harder to screen the background in the liquid drop model because ${\mathds{1}}_{\Omega}$ has a fixed height. Our idea is to split $D_{K}$ into many small boxes of size $\varepsilon$ and place in each full cube a smaller cube of side length $\rho^{1/3}\varepsilon$. Of course the main difficulty is to deal with the cubes intersecting the boundary of $D_{K}$ and this is where Proposition 10 becomes useful.
@@
Since the $\mu_{\alpha}$ are neutral, have no dipole and are contained in a set of diameter less that $C\varepsilon$, we have by  [AL20, Lemma 2.1]
@@ A.3. Upper bound for any regular set
Next we turn to the upper bound for an arbitrary sequence of sets $\Lambda_{n}\nearrow\mathbb{R}^{3}$ satisfying the conditions of Theorem 1. The idea is to again pack $\Lambda_{n}$ with many balls congruent to $B_{0},...,B_{K-1}$ for some fixed $K$. To this end, it is easier to start packing the large cube $C^{\prime}_{0}=|B_{K}|^{1/3}(-1/2,1/2)^{3}$ of the same volume as $B_{K}$ and then use this cube to pack $\Lambda_{n}$. By [LL72, Lemma 4.1] the cube $C^{\prime}_{0}$ can be packed by the same number $n_{K-j}$ of copies of the ball $B_{j}$, with the same estimate on the distance between the boundaries.
@@
$\mathcal{Z}^{\prime}_{b}\leqslant|B_{K}|^{-1}\left|\left\{x\in\Lambda_{n}\ :\ \mathrm{d}(x,\partial\Lambda_{n})\leqslant(\sqrt{3}+1)h_{K}\right\}\right|\leqslant C\ell_{n}^{2}h_{K}$
@@
$\mathcal{Z}^{\prime}_{i}=|B_{K}|^{-1}|\Lambda_{n}|+O(\ell_{n}^{2}h_{K}).$
@@
$\leqslant\#\mathcal{Z}^{\prime}_{i}\sum_{j=0}^{K-1}n_{K-j}E_{B_{j}}(\rho)+C\gamma^{K}|\Lambda_{n}|+C\ell_{n}^{2}h_{K}.$ (46)
@@
Figure 4. Idea of the proof of the thermodynamic limit for an arbitrary sequence $\Lambda_{n}\nearrow\mathbb{R}^{3}$, following [LL72, LN75]. To get an upper bound on $E_{\Lambda_{n}}(\rho)$ we pack the domain $\Lambda_{n}$ by many small balls, using a cubic tiling as guiding principle (only the largest balls of the type $B_{K-1}$ are displayed in the picture). To get a lower bound on $E_{\Lambda_{n}}(\rho)$ we pack the complement of $\Lambda_{n}$ in a large ball $B_{K_{n}}$ of comparable volume.
@@
Here $\mathcal{Z}^{\prime}_{i}$ denotes the sets of indices $z\in\mathbb{Z}^{3}$ such that $C^{\prime}_{z}$ is included in $B_{K_{n}}\setminus\Lambda_{n}$ at distance $h_{K}$ from its boundary.
@@
The next step is to handle the Coulomb interaction between the quadrupoles and $\Lambda_{n}$ in (47). We use that the Coulomb potential generated by a quadrupole decays like $1/|x|^{3}$ [AL20, Lemma 2.1], that is,
@@ A.5. The periodic problem
That the periodic model admits the same limit $e(\rho)$, as stated in (7), can be shown by following [LLS19, Section V]. Let us call
@@
To obtain the reverse bound, we pack $\Lambda_{L}$ with balls as explained in Subsection A.3 and in [LLS19]. We choose $\varepsilon\in L/\mathbb{N}$ to avoid having any partial cubes at the boundary of $\Lambda_{L}$ (and thus be consistent with the periodic boundary condition). Using that Newton’s theorem applies in the periodic case as well, we find the same bound as (46) on $E_{L}^{\rm per}(\rho)$ and this gives